Why wouldn't it be helpful to try and solve the following quadratic equation by taking square roots? 2x^2-2x-9=0
1
Expert's answer
2019-10-28T12:52:52-0400

"ax^2+bx+c=0" is a basic square equation. In our case "a=2; b=-2;c=-9" .

Solve the quadratic equation using the discriminant.

The discriminant is "D = b^2-4ac=(-2)^2-4*2*(-9)=76" .

"x_1=\cfrac{-b-\sqrt D}{2a}=\cfrac{2-\sqrt {76}}{4}; x_2=\cfrac{-b+\sqrt D}{2a}=\cfrac{2+\sqrt {76}}{4}" .

Answer: "x_1=\cfrac{2-\sqrt{76}}{4}; x_2=\cfrac{2+\sqrt{76}}{4}" .